# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

### v3

@@ Appendix C Data Cleaning Statistics
In this section, we provide detailed data cleaning statistics (Table 678 and 9) for high-resource, medium-resource, and low-resource languages. For the data cleaning statistics of very low-resource languages, please refer to the open-source data statistics we released.
@@ D.2 Pre-training and Evaluation Protocol
Given the limitations of computational resources, we refrain from full-parameter finetuning. Instead, we adopt Low-Rank Adaptation (LoRA; 53), a parameter-efficient fine-tuning technique that introduces trainable low-rank matrices into each transformer layer, substantially reducing the number of trainable parameters while maintaining competitive performance.
@@ D.3 Statistical Anomaly Detection
Appendix E Limitations
This work has the following limitations: (i) Although the proportion of high/medium/low resource languages in DCAD-2000 has greatly increased compare to existing multilingual datasets, a significant portion of the languages are still very low resource languages. Future work will explore to collect data for extremely low-resource languages through other modalities (e.g., images) through technologies like OCR. (ii) We evaluate the new data cleaning framework only on four classical anomaly detection algorithms; however, since the framework is algorithm-independent, it should also be effective with other anomaly detection algorithms. (iii) For language identification, we use GlotLID [21], a FastText-based model whose limitations in handling massive multilinguality have been discussed in previous works [54]. However, since the data cleaning pipeline is language-agnostic, other language identification models can also be employed.
Appendix F Ethics Statement

### v4

@@ Appendix C Data Cleaning Statistics
In this section, we provide detailed data cleaning statistics (Table 91011 and 12) for high-resource, medium-resource, and low-resource languages. For the data cleaning statistics of very low-resource languages, please refer to the open-source data statistics we released.
@@ D.2 Pre-training and Evaluation Protocol
Given the limitations of computational resources, we refrain from full-parameter finetuning. Instead, we adopt Low-Rank Adaptation (LoRA; 55), a parameter-efficient fine-tuning technique that introduces trainable low-rank matrices into each transformer layer, substantially reducing the number of trainable parameters while maintaining competitive performance.
@@ D.3 Statistical Anomaly Detection
Appendix E Manual Quality Evaluation of Cleaning Pipeline
To validate the effectiveness of our cleaning pipeline and to assess the residual noise and false positives, we conduct a manual quality evaluation of the retained and deleted documents. This evaluation was performed on five representative languages: English, Chinese, German, Japanese, and French. These languages are selected to ensure diverse linguistic coverage, and the evaluation will be extended in future work to include additional languages, particularly low-resource languages, where automatic filtering may be more challenging.
For each of the five languages, we randomly sampled 100 documents retained by our pipeline (i.e., documents that were kept) and 100 documents that were removed (i.e., deleted by our pipeline). The annotation process was conducted by one proficient annotator per language. The key goal of this annotation process was to estimate both the quality of the documents retained by the pipeline and the false positives in the deleted set. The quality evaluation provides insight into how well the cleaning pipeline separates high-quality content from noisy or irrelevant data. The documents were labeled with the following quality ratings:
  • Good: Documents that were coherent, meaningful, and of high quality.
  • Borderline: Documents that were understandable but flawed, including minor corruption, weak coherence, or other small issues.
  • Bad: Documents that were nonsensical, noisy, or semantically meaningless, such as machine translation errors, boilerplate content, spam, or mixed-language noise.
Table 4: Quality evaluation of retained and deleted documents across five languages.
Retained Documents (Kept by filter) Deleted Documents (Removed by filter)
Language Good Borderline Bad Residual Noise (Bad %) Language Good Borderline Bad False Positives (Good %)
English 86% 10% 4% 4% English 5% 14% 81% 5%
Chinese 82% 13% 5% 5% Chinese 6% 18% 76% 6%
German 84% 12% 4% 4% German 5% 16% 79% 5%
Japanese 81% 12% 7% 7% Japanese 6% 17% 77% 6%
French 84% 14% 2% 2% French 4% 15% 81% 4%
Avg 83.4% 12.2% 4.4% 4.4% Avg 5.2% 16% 78.8% 5.2%
Table 4 demonstrate that our cleaning pipeline effectively filters out low-quality content while preserving high-value data. Across all five languages, the proportion of retained documents rated as “Bad” (residual noise) averaged only 4.4%, indicating minimal contamination of retained documents by low-quality content. Similarly, the false positive rate (i.e., representing the proportion of high-quality documents mistakenly removed) was low, averaging 5.2%. The pipeline’s precision, defined as the proportion of retained documents classified as “Good” or “Borderline”, was 95.6%, while its recall, which measures the retention of “Good” documents, was 94.13%. These results demonstrate that the pipeline achieves both high precision and recall, effectively balancing the removal of noise with the preservation of valuable data. Overall, the findings validate the robustness of our unsupervised, anomaly-detection-based approach across multiple languages, with future work aimed at extending this evaluation to additional languages, particularly low-resource ones.
Table 5: Cost-Benefit Comparison of Filtering Methods (Per 1M Documents)
Metric Heuristic Filtering Anomaly Detection (DCAD)
Cleaning Time 10 minutes 12–15 minutes
Max Memory Usage (CPU) 58 GB 64 GB
Training Data Retained (%) 88% 77%
Avg Model Accuracy (Global MMLU subset) – LLaMA-3.2-1B 43.9% 48.6%
Accuracy Gain per CPU-hour +0.42% +0.60%
Accuracy Gain per 1% data lost +0.16% +0.32%
Appendix F Cost Benefit Analysis of Cleaning Strategies
To better evaluate the practical trade-offs between conventional heuristic filtering and our anomaly-based framework (DCAD), we conduct a controlled cost–benefit analysis on one million web documents under identical hardware conditions. As summarized in Table 5, the DCAD pipeline incurs only a minor computational overhead relative to the heuristic baseline (i.e., approximately two additional minutes of processing time and a 6 GB increase in peak memory usage). Although DCAD retains 11% fewer documents, it consistently yields superior downstream task performance (Section 5), highlighting its effectiveness in balancing data quality and computational efficiency.
Table 6: Ablation Study: Impact of Feature Subsets (Refer to Section 3.2.1)
Feature Subset Used Arabic Turkish
All 8 features (1–8) 0.21 0.27
w/o (8) Perplexity 0.20 0.25
w/o (7) LID score 0.16 0.21
w/o (6) Flagged word ratio 0.20 0.24
w/o (5) Stopword ratio 0.21 0.26
w/o (4) Special character ratio 0.20 0.26
w/o (3) Word repetition 0.18 0.24
w/o (2) Character repetition 0.19 0.25
w/o (1) Token count 0.20 0.24
Appendix G Feature Robustness Analysis
To evaluate the robustness of our anomaly detection framework with respect to feature design (Section 3.2.1), we conduct a one-feature-at-a-time ablation study. Given the combinatorial explosion of all possible subsets ($2^{8}-1=255$), we adopt a pragmatic protocol in which each feature is removed individually from the full 8-dimensional feature vector, and the cleaning process is repeated using the remaining seven features. We then fine-tune LLaMA-3.2-1B on each resulting filtered corpus and evaluate performance on FineTask–Arabic and FineTask–Turkish, following the same experimental setup as in Section 5.
As observed in Table 6, we have the following findings: (1) The full 8-feature configuration consistently outperforms all ablated variants, confirming that each feature contributes meaningfully to overall performance. (2) The Language Identification (LID) confidence score (Feature 7) is particularly critical: its removal results in a substantial accuracy drop, likely due to the presence of mixed or misidentified language content that adversely affects multilingual model quality. (3) Other features, such as repetition ratios and perplexity, provide modest gains individually; none are harmful or redundant when considered in isolation.
Appendix H Practical Choice of Anomaly Detector and Future Extensions
While modern deep anomaly detection methods, such as autoencoder-based reconstruction scoring [56] and contrastive outlier detection [57], have achieved strong performance in other domains, we deliberately adopt a classical algorithm, specifically Isolation Forest, in this work. This choice is motivated by three practical constraints inherent to large-scale multilingual corpus cleaning:
  • Lack of a clean reference distribution. Autoencoder-based methods assume access to a predominantly clean training set to learn a reliable reconstruction prior. In our weakly supervised scenario covering 2,282 languages without dependable clean subsets, this assumption is violated, making such models prone to degenerate reconstructions on noisy data.
  • Scalability across languages without supervision. Contrastive-learning-based outlier detection requires either labeled normal/abnormal pairs or implicitly curated positive anchors. Providing such supervision for thousands of languages would reintroduce the language-specific manual tuning that our language-agnostic pipeline explicitly avoids.
  • Resource efficiency and feature interpretability. Our framework relies on explicit, interpretable quality features (e.g., repetition ratio, perplexity, LID confidence) rather than opaque embedding-space distances. Classical anomaly detectors like Isolation Forest can operate directly on these CPU-computable features and scale to 46 TB of multilingual data without GPU dependency, making them well-suited for real-world data curation pipelines.
Nonetheless, extending DCAD to incorporate semantic embedding-based anomaly signals or lightweight deep novelty detection represents a promising direction for future work. We view our current feature-space approach as a foundational layer, onto which richer semantic detectors can be incrementally integrated once computational and language-coverage challenges are addressed.
Appendix I Ethics Statement